OpenStudy (anonymous):

integral of dy/sqrt(4-y^2)?

OpenStudy (freckles):

a trig sub will work just dandy here

OpenStudy (inkyvoyd):

you'll want to complete the square and what you're left with use trig sub

OpenStudy (anonymous):

\[\int\limits \frac{ dy }{ \sqrt{4-y ^{2}}}\]

OpenStudy (freckles):

Recall 1-sin^2(x)=cos^2(x) so 4-4sin^2(x)=4cos^2(x) so 4-(2sin(x))^2=4cos^2(x))

OpenStudy (freckles):

so it should be obvious what you want to replace y with

OpenStudy (anonymous):

i am not good with trig identities

OpenStudy (freckles):

that's fine i wrote the identity you needed

OpenStudy (freckles):

when is 4-(2sin(x))^2 the same as 4-y^2 when y= what?

OpenStudy (anonymous):

oh.. i see y=2sinx

OpenStudy (freckles):

so that is your sub

OpenStudy (anonymous):

ok thank you!

OpenStudy (freckles):

@inkyvoyd I'm interested in the partial fractions way

OpenStudy (freckles):

but how do you do it with the sqrt around the 4-y^2 part/

OpenStudy (inkyvoyd):

@freckles ... you don't. I am such an idiot.

OpenStudy (freckles):

oh

OpenStudy (anonymous):

wait freckles what happens after i sub in 2sinx

OpenStudy (anonymous):

what happens to sqrt

OpenStudy (freckles):

y=2sin(x) dy=?

OpenStudy (freckles):

And inky I don't believe you are an idiot. :p

OpenStudy (anonymous):

dy=2cosx

OpenStudy (inkyvoyd):

that's the problem @freckles ... that's how idiots get to the top of important positions :P

OpenStudy (freckles):

dy=2cos(x) dx

OpenStudy (freckles):

So replace dy with 2cos(x) dx replace y with 2sin(x)

OpenStudy (freckles):

and use the identity tht 4-4sin^2(x)=4cos^2(x)